Chapter 3 - Part 2 Problem 3.80 2 of7> If the bullet leaves the gun at a speed of 190 m/s, by how much will it miss the target? Constants 1 Periodic Table A hunter aims directly at a target (on the same level) 68.0 m away. Ay .27 Submit X Incorrect; Try Again; 4 attempts remaining Part B At what angle should the gun be aimed so the target will be hit? There are two possible answers, but only one is reasonable. ■循
Answer #1

Time taken by bullet, t = 68/190 = 0.36 sec

H = 0.5 x 9.8 x 0.36^2 = 0.63 m

B)

Range, R = v^2 sin2(theta)/g

Sin2(theta) = gR/v^2 = 9.8 x 68/190^2 = 0.018

Theta = 0.53 degree

Comment in case any doubt please rate my answer.....